OpenStudy (anonymous):

Please help write and solve this equation: Christopher's bill for lunch was $7.92. He had 2 burgers and a water. The water cost $.20 for the cup and the tax was $.54. How much was each burger?

OpenStudy (anonymous):

Total bill= 7.92. Cost of water is 0.20. Tax is .54. So the cost of the two burgers is 7.18. The cost for one burger is 7.18/2 (since he bought 2) is 3.59.

OpenStudy (anonymous):

That's great. But how would you write out the equation for that? Just the way you broke it down for me?

OpenStudy (anonymous):

\[7.92-(.20+.54)=n/2\]

OpenStudy (anonymous):

Is this right?

OpenStudy (anonymous):

7.92=.20+.54+2n. Total cost= tax+water+2*cost of one burger.